OpenStudy (anonymous):

factor the expression and use fundamental identities to simplify. 1- 2cos^2x+cos^4x

OpenStudy (lgbasallote):

you can rewrite it as \[1 - 2 (\cos x)^2 + (\cos x)^4\]

OpenStudy (lgbasallote):

now watch the magic.. let a = cos x 1 - 2a^2 + a^4 let b = a^2 1 - 2b + b^2

OpenStudy (lgbasallote):

are you following? or did i lose you?

OpenStudy (anonymous):

sorry sorry xD

OpenStudy (anonymous):

so why are we replacing it with variables? :o

OpenStudy (lgbasallote):

to make it look nicer and easier

OpenStudy (lgbasallote):

you can just use cos

OpenStudy (lgbasallote):

i substituted so you can see it clearly

OpenStudy (lgbasallote):

shall i continue?

OpenStudy (anonymous):

oh okay got it, go on

OpenStudy (anonymous):

haha sorry x)

OpenStudy (lgbasallote):

ahh good 1 - 2b + b^2 im going to rearrange it to make it look nicer b^2 - 2b + 1 now you can see this is factorable (b-1)(b-1) <--i hope you saw that

OpenStudy (lgbasallote):

now we sub b back to a b = a^2 so \[(a^2 - 1)(a^2 + 1)\] the first factor is still factorable so let's \[(a+1)(a-1)(a^2 + 1)\] now we sub a back to cosx a = cosx \[(\cos x + 1)(\cos x - 1)(\cos^2 x + 1)\]

OpenStudy (lgbasallote):

i hope you got my explanation. sorry for rushing. it's just that im about to leave lol

OpenStudy (anonymous):

awww :( i was just gonna say i got lost when you factored lol

OpenStudy (lgbasallote):

uhmm which part?

OpenStudy (lgbasallote):

b^2 - 2b + 1 = (b-1)(b-1)?

OpenStudy (lgbasallote):

oh no...my last post was wrong o.O do you get (b-1)(b-1) though?

OpenStudy (anonymous):

yeah that part

OpenStudy (lgbasallote):

it's a perfect square trinomial

OpenStudy (lgbasallote):

i dont know how to teach factoring.....it's a little complicated to explain...

OpenStudy (lgbasallote):

this is gonna be a little tricky...the next step would require a more advanced factoring...

OpenStudy (lgbasallote):

let me post the solution first then i'll tag someone to help explain it hehe...i need to go..

OpenStudy (anonymous):

okay sorry about that! but thanks for the help :)

OpenStudy (lgbasallote):

(b-1)(b-1) sub back to b = a^2 (a^2 - 1)(a^2 - 1) factor out a negative -(1 - a^2)(a^2 - 1) factor out another negative (1 - a^2)(1-a^2) sub a = cosx (1 - cos^2 x)( 1 - cos^2 x) (sin^2 x) (sin^2 x) sin^4 x

OpenStudy (lgbasallote):

@angela210793 could you help me explain the solution? i need to go :(

OpenStudy (lgbasallote):

sorry again @dorkkk